### v4

# Counterfactual Uncertainty Quantification of Factual Estimand of Efficacy from Before-and-After Treatment Repeated Measures Randomized Controlled Trials

Abstract
This article quantifies the uncertainty reduction achievable for counterfactual estimand, and cautions against potential bias when the estimand uses Digital Twins.
Posed by Neyman (1923a) who showed unbiased point estimation from designed factual experiments is possible, counterfactual uncertainty quantification (CUQ) remained an open challenge for about one hundred years. The $Rx:C$ counterfactual efficacy we focus on is the ideal estimand for comparing treatment $Rx$ with control $C$, the expected outcome differential if each patient received both $Rx$ and $C$.
Enabled by our new statistical modeling principle called ETZ, we show CUQ is achievable in Randomized Controlled Trials (RCTs) with Before-and-After Repeated Measures, common in many therapeutic areas. The CUQ we are able to achieve typically has lower variability than factual UQ.
We caution against using predictors with measurement error, which violates regression assumptions and can cause attenuation bias in estimating treatment effects. For traditional medicine and population-averaged targeted therapy, counterfactual point estimation remains unbiased. However, in both Real Human and Digital Twin approaches, estimating effects in subgroups may suffer attenuation bias.
Keywords: Variability decomposition, Randomized Controlled Trials, Digital Twins, Counterfactual Uncertainty Quantification, ETZ modelling principle, Before-and-After Repeated Measures.
1 Motivation and scope of application
When comparing a new treatment (which we denote by $Rx$) with a control (denoted by $C$), the ideal estimand is the counterfactual efficacy, the expected differential outcome between $Rx$ and $C$ if each patient were given both $Rx$ and $C$. Following Rubin (1978), causal treatment effect is defined as
“how the outcome of treatment compares to what would have happened to the same subjects under alternative treatment”
by the 2019 International Council for Harmonization guideline ICHE9(R1) on Estimands. From a patient’s point of view, indeed what matters is the personal treatment effect, the differential in outcomes if they switch from a control therapy $C$ to a new therapy $Rx$. We denote this as $Rx:C$ efficacy.
Most factual Randomized Controlled Trials (RCTs) are such that each patient is given either $Rx$ or $C$ but not both. However, counterfactual point estimation of counterfactual $Rx:C$ efficacy from factual studies is readily available, as random assignment of treatments to patients lets one prove that the point estimate of the differential $Rx$ and $C$ effects from modeling (average treatment effect (ATE)) a factual RCT is unbiased for the counterfactual $Rx:C$ efficacy.
The proof that designed experiments allow for unbiased counterfactual point estimation of the average treatment effect (ATE) from observable outcomes can be traced to Neyman (1923b), which has been edited and translated as Splawa-Neyman et al. (1990).
Importantly, how to quantify the uncertainty in estimating a counterfactual point estimate is a problem raised one hundred years ago but has remained unsolved. Neyman considered this problem “worthy of future study”. Neyman is unable to assess the variance of a counterfactual point estimate precisely, although he realized that the variance of a factual point estimate in standard practice is larger than what the variance of a counterfactual point estimate would be. See page 473 of Rubin (1990) which commented on Neyman (1923a).
Current uncertainty quantification of this point estimate is based on the variability of differential $Rx$ and $C$ effects between different patients, one treated with $Rx$ and the other with $C$, which intuitively is larger than the variability of differential effects if each patient were treated with both $Rx$ and $C$. With multiple components of variability, it is challenging to isolate and remove the person-to-person variability. Previous thinking was counterfactual uncertainty quantification requires studies with crossover designs and such which are rarely feasible. This article shows, actually, Counterfactual Uncertainty Quantification (which reduces the variance from factual uncertainty quantification) is achievable for a widely used kind of clinical trials that do not involve crossover designs.
Disease areas listed in Table 1 routinely conduct factual RCTs with each patient having Before-and-After treatment Repeated Measures (without crossing over). We call such studies factual BAtRM RCTs. We show, surprisingly, the uncertainty of $Rx:C$ efficacy point estimation from factual BAtRM RCTs can be quantified counterfactually. We do this by creating a new modeling principle, ETZ, to track individual patient outcomes in BAtRM RCTs and quantify their variability components. Disease areas in which factual BAtRM RCTs are routine affect half a billion patients worldwide, as shown in Table 1.
Therapeutic Area Number of Patients (in millions) Region
Crohn’s disease and ulcerative colitis 3 U.S.
Rheumatoid arthritis 54 U.S.
Schizophrenia 20 worldwide
Alzheimer’s disease 50 worldwide
Type 2 diabetes 400 worldwide
Table 1: Therapeutic areas with Before-and-After treatment Repeated Measures random controlled trials benefiting from the ETZ modeling principle.
Efficacy is primarily assessed at what is called the milestone visit. One can assess efficacy based on either the outcome measured at the milestone visit or change-from-baseline defined as the difference between the outcomes at the milestone visit and at baseline. The current practice in BAtRM RCTs uses change-from-baseline.
To gain insight, we propose a new statistical modeling principle called ETZ which separates variability of patient outcomes at the milestone visit into three components: variability in severity (stage of the disease that patients are initially at) $Z$, variability in trajectories $Traj$ of patients’ responses to treatments, and variability in the measurement errors $E$.
Our ETZ modeling principle in Section 4.5 shows using change-from-baseline instead of the milestone visit outcome reduces the uncertainty in efficacy assessment provided that the measurement error variability is not the main source of variability at the baseline visit (visit 1), a minimal requirement (for otherwise the outcome scale is of questionable usefulness). So the ETZ modeling principle validates current practice.
Further, without changing the (unbiased) treatment efficacy point estimation in current practice, our ETZ modeling principle allows what we call Counterfactual Uncertainty Quantification (CUQ) which calculates a (typically smaller) quantified variability of an idealized counterfactual estimand instead of the factual estimand. We illustrate CUQ using realistic examples of BAtRM RCTs.
A final insight from the ETZ modeling principle is that, since $Z$ is unobservable, it must be estimated. While baseline is often used, $Z$ can more generally be estimated from a set of biomarkers. However, when predictors are measured with error, standard regression analysis leads to attenuation of treatment effect estimates. Fortunately, we found that counterfactual point estimation remains unbiased for traditional medicine. For targeted therapies, it is also unbiased, as long as efficacy is defined as population-averaged (see Section 5.5). That said, all methods including our self-controlled human design and digital twins—should note that subgroup-level effect estimation may be biased if biomarkers are measured with error (see Section 5.4).
2 ETZ, a modeling principle
In practice, the differential effect between $Rx$ and $C$ is typically estimated from outcomes observed with some patients given $Rx$ and others given $C$.
While it is known that randomization allows unbiased estimation of a Counterfactual estimand from factual observations, Counterfactual Uncertainty Quantification (CUQ) of such an estimate has been largely thought as requiring crossover or N-of-1 studies since Neyman (1923a). We are happy to report that, in the broadly applicable setting of BAtRM RCTs, the ETZ modeling principle enables CUQ.
2.1 Conceptualizing an ideal counterfactual efficacy
Let $Trt_{i}$ be the random variable that assigns the $i^{th}$ patient to either $Rx$ or $C$. For the $i^{th}$ patient, let $Y_{i}(Rx)$ and $Y_{i}(C)$ represent potential outcomes if that patient is assigned to $Rx$ or $C$ respectively. Let $X$ represent relevant covariates in general. The mean effect of a treatment $Trt$ is the expected potential (possibly counterfactual) outcome if all individuals in the population had received the treatment $Trt$.
What is called an “ignorable” random assignment of patients to $Rx$ and $C$ is one such that, for each given $X=x$, the random assignment variable $Trt_{i}$ is independent of the potential outcomes $Y_{i}(Rx)$ and $Y_{i}(C)$.
For the $r^{th}$ patient assigned to $Rx$, let $Y^{Rx}_{r}$ represent the observed outcome, and for the $s^{th}$ patient assigned to $C$, let $Y^{C}_{s}$ represent the observed outcome.
Only $Y_{r}^{Rx}$ and $Y_{s}^{C}$ are observable, not $Y_{r}(C)$ or $Y_{s}(Rx)$. However, since $Trt_{i}$ is independent of potential outcomes under ignorable randomization, conditional on $Trt_{i}$ being either $Trt_{i}=Rx$ or $Trt_{i}=C$, potential outcomes and observable outcomes have the same expectations:
$\mathbbm{E}[Y_{i}(Rx)|X_{i}=x_{i}]$
$=$ $\mathbbm{E}\{\mathbbm{E}[Y_{i}(Trt_{i})|Trt_{i}=Rx,X_{i}=x_{i}]\}$
$=$ $\mathbbm{E}[Y^{Rx}_{i}|X_{i}=x_{i}],$
$\mathbbm{E}[Y_{i}(C)|X_{i}=x_{i}]$
$=$ $\mathbbm{E}\{\mathbbm{E}[Y_{i}(Trt_{i})|Trt_{i}=C,X_{i}=x_{i}]\}$
$=$ $\mathbbm{E}[Y^{C}_{i}|X_{i}=x_{i}].$
Assuming the distribution of $X$ is the same for $Rx$ and $C$, taking expectation over the distribution of $X$, if efficacy $\tau$ is defined as the expected difference of potential outcomes, then
$\tau$ $=$ $\mathbbm{E}[Y(Rx)-Y(C)]$ (1)
$=$ $\mathbbm{E}_{X}[\mathbbm{E}(Y(Rx)-Y(C)|X=x)]$
$=$ $\mathbbm{E}_{X}[\mathbbm{E}(Y(Rx)|X=x)-\mathbbm{E}(Y(C)|X=x)]$
$=$ $\mathbbm{E}_{X}[\mathbbm{E}(Y^{Rx}|X=x)-\mathbbm{E}(Y^{C}|X=x)]$ (2)
$=$ $\mathbbm{E}_{X}[\mathbbm{E}(Y^{Rx}|X=x)]-\mathbbm{E}_{X}[\mathbbm{E}(Y^{C}|X=x)]$ (3)
$=$ $\mathbbm{E}[Y^{Rx}]-\mathbbm{E}[Y^{C}]$ (4)
so this efficacy in the population space with infinitely many patients can be assessed based on observable outcomes.
3 Conceptualizing an ideal personal $Z$ covariate
The ETZ modeling principle conceptualizes each patient as having their own random intercept $Z$ before any treatment is given. Trajectory, denoted by $Traj$, is the random response from each patient to the treatment $Trt$ given. Within each $Trt$, trajectory $Traj$ is randomly distributed around a profile of long run average trajectories (without a particular shape specified). ETZ is a counterfactual modeling principle, with observable outcomes given the factual treatment, and potential outcomes given the counterfactual treatment.
Compared to the usual baseline predictor, the $Z$ component has two unique features: it is unobservable and without measurement error. As visualized in Figure 1, neither the intercept $Z$ nor the trajectory curves are directly observable, because there are measurement errors. Observed and potential outcomes are represented as dots and squares respectively in Figure 1.
We call ETZ a counterfactual modeling principle because $Z$ is not observable. Each individual is assigned a scalar $Z$ for their covariate $X$, on the $Y$ outcome scale. The ETZ modeling principle uses $Z$ to model biological variability among patients, conceptualizing initial intercept value $Z$ as predictive of outcome.
The ETZ modeling principle reduces the quantified uncertainty of estimated counterfactual efficacy at the population level by first using each patient as their own control (as in current practice), then further reduces it by calculating counterfactually under the ETZ modeling principle.
The ETZ modeling principle cannot be applied directly since $Z$ is unobservable. Model fitting has to be done factually, so (for proof-of-concept) we model the data as Mixed Model Repeated Measures (MMRM) using baseline measurement as a surrogate to $Z$. We then make a 1-to-1 transformation between the variability components in MMRM and ETZ. We call this process Real Human Counterfactual Estimation (RHCE).
3.1 A factual Digital Twins model
Instead of conceptualizing patients as having different (unobservable) initial values $Z$, one can also account for patient variability by their having different observable biomarker values $(b_{1},\ldots,b_{p})$ such as age, baseline or other measurement of disease severity, sex, etc. An algorithm $g$ can be trained to predict an $Rx$ patient’s digital twin’s $C$ outcome and a $C$ patient’s digital twin’s $Rx$ outcome based on $(b_{1},\ldots,b_{p})$. In what we call Digital Twin Counterfactual Estimation (DTCE), after calculating the difference between each patient’s factual and its digital twin’s counterfactual outcome, population $Rx:C$ efficacy is estimated by averaging them across the patients. See National Academies of Sciences, Engineering, and Medicine (2024), for example, for such thinking. DTCE can also potentially reduce uncertainty of estimated efficacy.
Beware of covariates with measurement error
It is important to be cognizant of covariates measured with error (e.g, baseline or other measures of disease severity). Predicting outcomes based on such covariates violates the standard regression modeling assumption that predictors are measured without error.
Inference from models with measurement error in predictors, called errors-in-variables models, will have attenuation bias. Therefore, in Section 5.5, we explain the care that both RHCE and DTCE need to take to avoid this bias. To place both RHCE and DTCE in the same framework, we take the algorithm $g$ as being trained to create a digital patient value $B=g(b_{1},\ldots,b_{p})$ similar to baseline, which is used as one predictor for patient’s $Y$ outcome, given either $Rx$ or $C$.
3.2 ETZ notations
Index each of the $m$ visits of a repeated measures study by $v$, $v=1,\ldots,m$. Denote by $Y^{Trt[v]}$ the outcome of a patient given treatment $Trt$ (either $Rx\mbox{ or }C$) measured at visit $v$. Values of profiles for $Rx$ and $C$ at times of visits are the fixed effects in the MMRM model (6). The ETZ modeling principle does not assume a functional form for the profiles, neither does MMRM.
To be more specific, conceptualize $Y^{Trt[v]}$ as having either two or three random components. At visit 1 ($v=1$), before any treatment is given, $Y^{Trt[1]}$ has two random components, a component $Z$ which is that patient’s intercept (true baseline measurement), and a measurement error component $E$ which is the difference between the true and the observed baseline. Thereafter, $Y^{Trt[v]}$ has three random components, a $Z$ which is that patient’s intercept (true baseline measurement), plus a random trajectory $Traj$ component (around its mean profile $\mu$) representing that patient’s response to the treatment given up to the time of visit $v$ ($v>1$), and a measurement error component $E$ which is the difference between the measured outcome and that patient’s true trajectory at each visit $v$. These components can be seen in Figure 1. It is assumed that measurement error $E$ is independent of $Z$ and $Traj$, and i.i.d. over the visits, across all the patients.
Figure 1: Components of the ETZ modeling principle for a RCT with $m=4$ visits at Time = 0, 1, 2, and 3. Black triangle is the baseline of patient 1 before any treatment is taken. Blue curve is the trajectory of patient 1 who is assigned to $C$, and the blue dots are the observed $Y$. Red curve represents patient 1’s potential trajectory had they been assigned to $Rx$, and the red squares are the potential $Y$ (with last value missing). Grey curve is the trajectory of patient 2 who is assigned to $Rx$. Measurement errors are differences between the outcomes and their trajectories.
3.3 Independence as a starting point
Our discussion of the ETZ modeling principle starts in the setting where the random intercept $Z$ and the random trajectory $Traj$ are independent within each treatment arm and goes toward where the direction of dependence between $Z$ and $Traj$ is known (with consequence to be discussed in Section 4.4).
Independence means that, the true stage of the disease that a patient is at before they receives either $Rx$ or $C$ does not inform on whether that patient’s response trajectory is going to be above or below the profile $\mu$ (long run average of the trajectories). By the Monotone Class Theorem, it is sufficient to check whether $P\{B|A\}=P\{B\}$ for events $A$ and $B$ of the form $B=\{improve>b\}$ and $A=\{severity>a\}$. That is, within each treatment arm, whether knowledge of severity being greater than 25 (say) informs on whether a patient would improve by more than 10 points (say). (Note that independence between the random effects $Z$ and $Traj$ within each treatment arm is unrelated to whether the fixed effect response profiles are parallel or not.)
If patient-level data is available, then one way to check independence is to see how baseline measurement is predictive of change-from-baseline. We take an AbbVie Alzheimer’s disease (AD) data as an example. Schnell et al. (2017) described an Alzheimer Disease study from AbbVie which compared three doses of a new treatment (doses 1, 2, and 3) with a negative control (dose 0, a placebo). There was an active control (dose 4) in the study as well. Outcome measure was ADAS-cog11, and duration of this Phase 2 study was 24 weeks. For the AbbVie AD data, the R-square regressing change-from-baseline (negative of the “improve” variable in the data set) on baseline (variable “severity” in the data set) allowing different profiles for the placebo and the 4 doses is 0.17. So it appears severity is not very predictive of trajectory.
A subtle concern is, regression modeling assumes predictors are measured without error, what if some of them have measurement error? Fortunately, as will be explained in Section 5.5, as long as efficacy is defined as “on average”, estimated $Rx:C$ efficacy is unbiased. However, that Section 5.4 also shows, in predicting efficacy in subgroups or at particular biomarker values, ignoring measurement error in predictors will lead to bias.
3.4 Properties of a counterfactual model
If the outcome of $i^{th}$ patient given treatment $Trt_{i}$, $Y_{i}(Trt_{i})$, is thought to only depend on the treatment $Trt_{i}$ and the covariate $X_{i}$, then the model is said to be deterministic counterfactual.
In real life, outcome measure on a patient is non-deterministic. Each potential measurement on each patient, if replicated, would be somewhat different. Reason for such variability includes the patient having a good day or a bad day (if they are asked to perform some tasks), rater variability, measurement error in analytic chemistry. If, in addition to the treatment effect $Trt_{i}$ and the covariate effect $X_{i}$, there is random measurement variability affecting $Y_{i}(Trt_{i})$, then the model is said to be non-deterministic counterfactual. In standard statistical notations, we write such a model as
$Y_{i}(Trt_{i})=h(Trt_{i},X_{i})+\epsilon_{i}$ (5)
where $h(Trt_{i},X_{i})$ represents treatment and covariate effects and $\epsilon_{i}$ represents measurement variability with $E(\epsilon_{i})=0$, assuming $\epsilon_{i}$ is independent of $(Trt_{i},X_{i})$. Since $E(\epsilon_{i})=0$, the derivation leading to the conclusion (4) that $\tau=E(Y^{Rx})-E(Y^{C})$ remains valid.
The counterfactual model represents an ideal situation, with the following properties.
Unbiased efficacy
With each patient receiving both $Rx$ and $C$, the $X$ values are perfectly matched between the $Rx$ and $C$ outcomes. So $E[Y(Rx)-Y(C)]$ is not biased by patients having high $X$ values systematically tending to receive $Rx$ while patients with low $X$ values tending to receive $C$ (say).
Reduced variability
Intuitively, $Y(Rx)-Y(C)$ with both $Rx$ and $C$ given to the same patient has less variability than $Y^{Rx}-Y^{C}$ giving $Rx$ to one patient and giving $C$ to a different patient.
The ETZ modeling principle we propose in this article leads to model (5) that separates measurement variability $\epsilon_{i}$ (the “E”) from patient-to-patient variability in $h(Trt_{i},X_{i})$ and crucially further separates the latter to before treatment (the “Z”) and after treatment (the “T”) variability.
3.5 BAtRM for counterfactual inference
To achieve counterfactual inference by giving both $Rx$ and $C$ to each patient is unethical for diseases that cause mortality (e.g., cancer) or irreversible mobility (e.g., arthritis).
Consider instead a factual RCT in which each patient first gets a baseline measurement $Y^{[1]}$ before being randomized to receive either $Rx$ or $C$, and then another measurement $Y^{[m]}$ at the milestone time point when efficacy in the primary endpoint is assessed. We call such RCTs Before-and-After treatment Repeated Measures (BAtRM) studies.
Therapeutic areas with such RCTs include type 2 diabetes mellitus (T2DM), neurodegenerative diseases such as Alzheimer’s disease (AD), psychiatric diseases such as schizophrenia, immunological diseases including rheumatoid arthritis, ulcerative colitis, Crohn’s disease, and psoriasis, afflicting hundreds of millions of patients worldwide in total.
Counterfactual inference includes two aspects: counterfactual (point) estimation and counterfactual uncertainty quantification. Surprisingly, counterfactual inference is possible from factual baseline and milestone RCTs, following a 3-step process. The first step achieves counterfactual estimation by Gauss-Markov theorem using factual sample, while the rest two steps explains the variance reduced by counterfactual uncertainty quantification from factual case.
Balancing
For a finite factual sample (which will not have perfectly balanced, let alone perfectly matched, $X$ values between $Rx$ and $C$), Gauss-Markov covariate adjustment ensures unbiased counterfactual point estimation by mapping the sample space with imbalance to an ideal balanced population space, as we demonstrate in Section 5.4.
Baselining
In RCTs with baseline and milestone measurements, variability incurred by giving $Rx$ to one set of patients and $C$ to the other patients can largely be removed by measuring outcome as change-from-baseline ($Y^{[m]}-Y^{[1]}$), as we show in Section 4.5.
Self-controlling
Merely providing an unbiased point estimate of efficacy with reduced variability is insufficient for statistical inference, an accurate assessment of the variability of the estimate is needed as well. Section 4.4 shows how to transform the following variabilities from a factual model
$\sigma^{B}$
Standard deviation of baseline measurements
$\sigma^{Mile}$
Standard deviation of outcome at the milestone time point
$\sigma^{Change}$
Standard deviation of change-from-baseline
to the three variabilities in the ETZ modeling principle, from which variability of our counterfactual estimand can be calculated.
ETZ is a modeling principle applicable to any BAtRM data structure, not restricted to any specific analysis method. How to estimate $\sigma^{B}$, $\sigma^{Mile}$, and $\sigma^{Change}$ from factual data is not unique. For proof of concept, we show how they can be obtained from widely used Mixed Model Repeated Measures (MMRM) analyses introduced in Section 4.1, and then transformed 1-to-1 to variabilities in the ETZ modeling principle in Section 4.4.
4 Factual modeling and counterfactual uncertainty quantification
The setting considered by MMRM is a longitudinal Randomized Controlled Trial in which there are $m$ planned visits (visit 1, $\ldots,m$) for all the patients. The times of these visits after randomization are the same for all the patients, and visit $m$ is the milestone time point at which efficacy in the primary endpoint is assessed. Denote by $Y^{[v]}_{i}$ the measurement on patient $i$ during visit $v$.
4.1 Mixed Model Repeated Measures (MMRM)
The so-called Mixed Model Repeated Measures (MMRM) representation is
$Y_{i}^{[v]}=\mu+\alpha^{[v]}+\beta^{Trt}+(\alpha\beta)^{Trt[v]}+b_{i}^{Trt}+w_{i}^{Trt[v]},$ (6)
$~{}v=1,\ldots,m.$
Fixed effects are represented by $\mu+\alpha^{[v]}+\beta^{Trt}+(\alpha\beta)^{Trt[v]}$, the mean response under treatment $Trt$ at time $v$, where $\alpha^{[v]}$ is the visit (time) effect, $\beta^{Trt}$ is the treatment effect, and $(\alpha\beta)^{Trt[v]}$ is the treatment×$\times$×visit interaction. Random between-subject effect for the $i^{th}$ patient assigned to treatment $Trt$ is $b_{i}^{Trt}$, while random within-subject effect is represented by $w_{i}^{Trt[v]}$.
[…]
$\sigma^{B}$ $=\sqrt{Var(Y_{i}^{[1]})}$
$\sigma^{Mile}$ $=\sqrt{Var(Y_{i}^{[m]})}$ (7)
$\sigma^{Change}$ $=\sqrt{Var(Y_{i}^{[m]})+Var(Y_{i}^{[1]})-2Cov(Y_{i}^{[m]},Y_{i}^{[1]})}$
[…]
The standard MMRM analysis from which we obtain these estimates assumes the missing data mechanism is MAR (Missing At Random). Strategies for dealing with other missing mechanisms are given in Chapter 19 of Mallinckrodt et al. (2019).
4.2 Comparison between MMRM and ETZ modeling
The ETZ modeling principle conceptualizes BAtRM clinical data biologically, with each of $E$, $Traj$, and $Z$ medically interpretable. ETZ separates the between subject effect $b_{i}^{Trt}$ in the MMRM model (6) into two parts: a visit 1 before-treatment $Z$ part, and a treatment-dependent $Traj$ part starting with visit 2. ETZ lets us articulate control of variability. Specifically, while intercept variability $Var(Z)$ can be reduced by narrowing the patient entry criterion, trajectory variability $Var(Traj)$ can be reduced by decreasing drug target heterogeneity among patients. Variability in measurement error $Var(E)$ can be reduced by using better trained raters, for example.
Mixed Model Repeated Measures (MMRM) is not necessary for the ETZ modeling principle but only one of possible methods which provides ETZ modeling with the information needed. We use MMRM for proof-of-concept as it is widely used in drug development currently. However, other methods such as fitting a random coefficient model (RCM) when appropriate can be used alternatively.
4.3 Plausibility
Intuitively, variability at the first and milestone visits and of Change may contain all the information we need to separate the $E$, $Traj$, and $Z$ variability components. Baseline measurements at the beginning of the study (visit 1) have the random intercept $Z$ and measurement error $E$ components but no random trajectory component $Traj$ since patients have not been treated yet. Change, or change-from-baseline, measurement at the milestone visit minus measurement at the first visit, has the random trajectory $Traj$ and the measurement error $E$ components (two copies of $E$ in fact) but not the random intercept component since that has been subtracted off. Measurements at the end of the study (visit $m$) have all three random components. So, assuming variability of measurement errors is constant across the visits, it seems plausible that $Var(Y^{[1]}_{i})$, $Var(Y^{[m]}_{i})$, and $Var(Y^{[\mathrm{change}]}_{i})$ are sufficient for us to recover the $E$, $Traj$, and $Z$ variability components.
4.4 Components of the ETZ decomposition
Our discussion starts without assuming variability of measurement error at visit $m$ is the same as variability of measurement error at visit $1$, or that the random intercept $Z$ is independent of the random trajectory $Traj$. As the decomposition proceeds, these two assumptions will be explicitly added as needed.
With such a conceptualization, outcomes of patient $r$ given $Rx$ and patient $s$ given $C$ at visit 1 and visit $m$ are represented as
$Y^{Rx[1]}_{r}$ $=$ $Z_{r}+E^{[1]}_{r}$ (8)
$Y^{Rx[m]}_{r}$ $=$ $Z_{r}+Traj^{Rx[m]}_{r}+E^{[m]}_{r}$ (9)
$Y^{C[1]}_{s}$ $=$ $Z_{s}+E^{[1]}_{s}$
$Y^{C[m]}_{s}$ $=$ $Z_{s}+Traj^{C[m]}_{s}+E^{[m]}_{s}.$
Implicitly, the fixed effects in this model are
$\mathbbm{E}(Z_{r})=\alpha_{Rx}=\mathbbm{E}(Z_{s})$ $=$ $\alpha_{C}$ (10)
$\mathbbm{E}(Traj^{Rx[m]}_{r})$ $=$ $\mu^{Rx[m]}_{r}$ (11)
$\mathbbm{E}(Traj^{C[m]}_{s})$ $=$ $\mu^{C[m]}_{s}$ (12)
with $\mathbbm{E}(Z_{r})=\mathbbm{E}(Z_{s})$ because patients are randomly assigned to $Rx$ and $C$.
Assuming variabilities are the same under $Rx$ and $C$, for patient $r$ given treatment $Trt$ which can be $Rx$ or $C$,
$Var(Y^{Trt[1]}_{r})$ $=$ $Var(Z_{r})+Var(E^{[1]}_{r})$ (13)
$Var(Y^{Trt[m]}_{r})$ $=$ $Var(Z_{r})+Var(E^{[m]}_{r})$
$+2~{}Cov(Z_{r},Traj^{Trt[m]}_{r})$
$+Var(Traj^{Trt[m]}_{r}),$
where $Z_{r}$ is the random intercept, $E^{[1]}_{r}$ is the random visit 1 measurement error, $E^{[m]}_{r}$ is the random visit $m$ measurement error and $Traj^{Trt[m]}_{r}$ is the random trajectory.
Note
$Var(Z_{r})+Cov(Z_{r},Traj^{Trt[m]}_{r})$ $=$ $Cov(Y^{Trt[m]}_{r},Y^{Trt[1]}_{r})$
and, at this point, we have five unknown ETZ variability components:
$Var(Z_{r}),Cov(Z_{r},Traj^{Trt[m]}_{r}),Var(Traj^{Trt[m]}_{r}),Var(E^{[1]}_{r}),$
$\mathrm{~{}and~{}}Var(E^{[m]}_{r}).$
Change, or change-from-baseline, is measured by subtracting the baseline (visit 1) measurement from the milestone (visit $m$) measurement, so it does not have a random intercept component. Change of patient $r$ given treatment $i$ is then
$Y^{Trt[\mathrm{change}]}_{r}=\left[Traj^{Trt[m]}_{r}\right]+\left[E^{[m]}_{r}-E^{[1]}_{r}\right]$
so
$Var(Y^{Trt[\mathrm{change}]}_{r})$ $=$ $Var(E^{[1]}_{r})$
$+Var(Traj^{Trt[m]}_{r})$
$+Var(E^{[m]}_{r})$
which leads to
$~{}Var(Z_{r})+Cov(Z_{r},Traj^{Trt[m]}_{r})$
$=\dfrac{Var(Y^{Trt[m]}_{r})+Var(Y^{Trt[1]}_{r})-Var(Y^{Trt[\mathrm{change}]}_{r})}{2}$ (16)
$=Cov(Y^{Trt[m]}_{r},Y^{Trt[1]}_{r})$ (17)
with the number of unknown ETZ variability components remaining five. (Note that since estimates of the three variances in (16) are routinely included as summary statistics in reporting RCT outcomes, $Cov(Y^{Trt[m]}_{r},Y^{Trt[1]}_{r})$ in (17) can be estimated without access to patient level data.) Making the assumption $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$, reducing the number of unknowns to four, gives
$Var(Traj^{Trt[m]}_{r})+2Cov(Z_{r},Traj^{Trt[m]}_{r})$ (18)
$=$ $Var(Y^{Trt[m]}_{r})-Var(Y^{Trt[1]}_{r}).$
If we further make the assumption that $Z$ and $Traj$ are independent, so that $Cov(Z_{r},Traj^{Trt[m]}_{r})=0$, reducing the number of unknown variability components to three, then we can estimate $Var(Z_{r})$ from (16), and estimate $Var(Traj^{Trt[m]}_{r})$ from (18). In turn, we can estimate variance of measurement error as
$Var(E^{[1]}_{r})$ $=$ $Var(Y^{Trt[1]}_{r})-Var(Z_{r})$
$=$ $Var(E^{[m]}_{r}).$
Figure 2 examines limitations of current ETZ modeling with the independence assumption. One potential issue is, if $Z$ and $Traj$ are in fact positively correlated, then from (17), (18), and (4.4), estimates of $Var(Z_{r})$ and $Var(Traj^{Trt[m]}_{r})$ would be higher than they ought to be, while the estimate for $Var(E^{[1]}_{r})=Var(E^{[m]}_{r})$ would be lower than it ought to be. Then we might miss the signal that $Var(E)>Var(Z)$, in which case study result is unreliable because within-patient measurement variability dominates between-patient variability. The right panel of Figure 2 shows how one can make a plot to assess how large the correlation needs to be for this signal to be missed.
Figure 2: Ratio SD($E$)/SD($Z$) of standard deviations across different clinical endpoints (ADAS-Cog, iADL, and MMSE) as functions of the correlation between $Z$ and $Traj$, using EXPEDITION3 as an example. The left panel shows if $Z$ and $Traj$ are sufficiently negatively correlated, then ETZ modeling can result in $Var(E)<0$ clearly indicating the independence assumption does not hold. The right panel shows the extent to which positive correlation between $Z$ and $Traj$ can cause ETZ modeling with the independence assumption to miss the signal that $Var(E)>Var(Z)$. Purple dashed line indicates when SD ($E$) reaches 80% of SD ($Z$), while orange dashed line indicates SD ($E$) = SD ($Z$).
Another potential issue is, if $Z$ and $Traj$ are sufficiently negatively correlated, then as shown in the left panel of Figure 2, ETZ modeling can result in $Var(E)<0$ which is nonsensical. But we do not expect negative correlation to be common in practice, as that corresponds to patients at late stage of the disease having more favorable trajectories.
4.5 Uncertainty reduced by counterfactual uncertainty quantification
We now show exactly how Counterfactual Uncertainty Quantification (CUQ), the goal of our research, is achieved by the ETZ modeling principle.
Narrowing patient entry criterion is ineffective
There is the thinking that narrowing the patient entry criterion makes patients more homogeneous and thereby makes clinical trials more likely to succeed. This turns out to not be the case if treatment effects are measured in terms of change-from-baseline, because (4.4) in the ETZ decomposition shows SD(Change) does not involve patients’ baseline measurements $Z$.
Trulicity® (compound name dulaglutide) is a once-weekly injectable prescription medicine to improve blood sugar (glucose) in adults with type 2 diabetes (mellitus), as measured by control of haemoglobin A1c (HbA1c). Biologic License Application 125469 for Trulicity/dulaglutide had five studies. Patient entry criteria, SD(Basline), and SD(Change) for studies GBDA and GBDC, reported on pages 6 and 7 in FDA center for drug evaluation and research (2014b) and on pages 58 and 61 of FDA center for drug evaluation and research (2014a), are reproduced in Tables 2 and 3.
HbA1c Dulaglutide + Metformin + Pioglitazone Placebo Exenatide
$n$ 279 141 276
SD(Baseline) 1.3 1.3 1.3
SD(Change) 0.98 0.97 1.02
Table 2: GBDA sub-study: Dulaglutide + Metformin + Pioglitazone vs. Placebo and Exenatide with baseline HbA1c $\in[7,11]$
HbA1c Dulaglutide Metformin
$n$ 269 268
SD(Baseline) 0.9 0.8
SD(Change) 1.00 0.98
Table 3: GBDC sub-study: Dulaglutide vs. Metformin with baseline HbA1c $\in[6.5,9.5]$
As can be seen, narrowing the HbA1c baseline entry criterion from [7, 11] in GBDA to [6.5, 9.5] in GBDC decreases SD(baseline) of HbA1c from 1.3 to within [0.8, 0.9], but hardly affects SD(Change), which remains steady within [0.97, 1.02] for both studies. So decreasing variability of intercept $Z$ may not reduce CUQ.
Baselining significantly reduces uncertainty
ETZ pinpoints when baselining effectively reduces uncertainty.
We show below that baselining effectively replaces the patient-to-patient variability $Var(Z)$ by the measurement error variability $Var(E)$ and, provided that $Var(Z)>Var(E)$, baselining is beneficial.
Whichever treatment patient $i$ is assigned to, expressions for $Var(Y_{i}^{[m]})$ and $Var(Y_{i}^{[change]})$ are
$Var(Y_{i}^{[m]})$ $=$ $Var(Z_{i})+Var(Traj_{i})+Var(E_{i}^{[m]})$ (20)
$Var(Y_{i}^{[1]})$ $=$ $Var(Z_{i})+Var(E_{i}^{[1]})$ (21)
$Var(Y_{i}^{[change]})$ $=$ $Var(Y_{i}^{[m]}-Y_{i}^{[1]})$ (22)
$=$ $Var(Z_{i}+Traj_{i}+E_{i}^{[m]}-Z_{i}-E_{i}^{[1]})$
$=$ $Var(Traj_{i}+E_{i}^{[m]}-E_{i}^{[1]})$
$=$ $Var(Traj)+2\times Var(E).$
with (22) showing the expression for the variance of change-from-baseline. So the difference between $Var(Y_{i}^{[m]})$ and $Var(Y_{i}^{[change]})$ is
$Var(Y_{i}^{[m]})-Var(Y_{i}^{[change]})$ (23)
$=$ $Var(Z)-Var(E)$
and $Var(Y_{i}^{[change]})<Var(Y_{i}^{[m]})$ when $Var(E)<Var(Z)$. (Note that if measurement variability is larger than patient-to-patient variability, $Var(E)>Var(Z)$, then it is hopeless for that clinical trial to be informative and a different outcome measure should be chosen.) Figure 3 in Section 4.6 shows that baselining can reduce total variability by 73% in an example data set.
Self-controlling CUQ further reduces uncertainty
Beyond baselining, self-controlling CUQ further reduces uncertainty by two copies of the measurement error variance, using ETZ to quantify uncertainty counterfactually instead of factually.
Factual uncertainty quantification considers a randomly selected pair of patients with patient $r$ given $Rx$ and patient $s$ given $C$. With $Y_{r}^{Rx[change]}$ and $Y_{s}^{C[change]}$ denoting their change-from-baseline outcomes, factual efficacy $Y^{f}$ is estimated by
$Y^{f}=Y_{r}^{Rx[change]}-Y_{s}^{C[change]}.$
Self-controlling uncertainty quantification considers the difference between potential change-from-baseline of the same patient $i$ receiving both $Rx$ and $C$. With $Y_{i}^{[change]}(Rx)$ and $Y_{i}^{[change]}(C)$ denoting the potential change-from-baseline outcomes for patient $i$, truly counterfactual efficacy $Y^{cf}$ is estimated by
$Y^{cf}=Y_{i}^{[change]}(Rx)-Y_{i}^{[change]}(C).$
ETZ assumes $Var(E)$ remains constant and $Var(Traj)=Var(Traj^{Rx})=Var(Traj^{C})$. Moreover, $E$ and $Traj$ are independent because how patients respond to treatment is irrelevant to measurement error.
Hence, factual variability is
$Var(Y^{f})$ $=$ $Var(Y_{r}^{Rx[change]}-Y_{s}^{C[change]})$ (24)
$=$ $Var(Traj_{r}^{Rx}+E_{r}^{[1]}+E_{r}^{Rx[m]}$
$-Traj_{s}^{C}-E_{s}^{[1]}-E_{s}^{C[m]})$
$=$ $2\times Var(Traj)+4\times Var(E),$
where $Traj_{r}^{Rx}$ and $Traj_{s}^{C}$ are independent because patients $r$ and $s$ are different patients. On the other hand, counterfactual variability is
$Var(Y^{cf})$ $=$ $Var(Y_{i}^{Rx[change]}-Y_{i}^{C[change]})$ (25)
$=$ $Var(Traj_{i}^{Rx}+E_{i}^{[1]}+E_{i}^{Rx[m]}$
$-Traj_{i}^{C}-E_{i}^{[1]}-E_{i}^{C[m]})$
$=$ $Var(Traj_{i}^{Rx}+E_{i}^{Rx[m]}-Traj_{i}^{C}-E_{i}^{C[m]})$
$=$ $2\times Var(Traj)+2\times Var(E)$
$-2\times Cov(Traj_{i}^{Rx},Traj_{i}^{C}).$
Thus,
$Var(Y^{f})-Var(Y^{cf})$ (26)
$=$ $2\times Var(E)+2\times Cov(Traj_{i}^{Rx},Traj_{i}^{C}).$
where $Cov(Traj_{i}^{Rx},Traj_{i}^{C})$ is non-negative because trajectories from the same patient $i$ to $Rx$ and $C$ would be independent or positively correlated as explained below.
Biologically, if $C$ is a placebo while $Rx$ is a targeted therapy, then one can expect the trajectories to be independent. If both $Rx$ and $C$ are targeted therapies but with entirely different targets, for example $Rx$ is GLP-1 while $C$ is metformin in treating T2DM patients, then one would expect the trajectories to be independent.
Trajectories can expect to be positively correlated when $Rx$ and $C$ have similar mechanism of action or are upstream and downstream of the same cascade or pathway. For example, $Rx$ and $C$ for treating schizophrenia may both be dopamine receptor antagonists. For treatment of Alzheimer’s disease, $C$ might target $\beta$-amyloid plaque (upstream) while $Rx$ targets tau-tangle (downstream).
So self-controlling removes at least two copies of the measurement error variance, and our claim of CUQ reducing $2\times Var(E)$ may even be conservative. Figure 3 in Section 4.6 shows that self-controlling CUQ can reduce an additional 17% of uncertainty beyond baselining in a real example.
4.6 Real data examples of CUQ
In this section, we show the proportion of variance reduced by counterfactual uncertainty quantification using two real clinical datasets, the AbbVie dataset and the EXPEDITION3 dataset.
Schnell et al. (2017) described an Alzheimer Disease study from AbbVie which compared three doses of a new treatment (doses 1, 2, and 3) with a negative control (dose 0, a placebo). The AbbVie data set is described in Section 3.3.
Honig et al. (2018) reported on EXPEDITION3, a double-blind, placebo-controlled, Phase 3 study testing solanezumab on patients with mild dementia due to Alzheimer’s disease. Among a total of 2129 patients enrolled, 1057 were assigned to receive solanezumab while 1072 were assigned to the placebo. Score at the initial (Week 0) visit before randomization was considered baseline. Thereafter outcomes were measured at weeks 12, 28, 40, 52, 64, and 80. Change-from-baseline is the difference between milestone outcome (at Week 80) and baseline (at Week 0). In EXPEDITION 3, the primary endpoint was the 14-item cognitive subscale of the Alzheimer’s Disease Assessment Scale (ADAS-cog14), while the key secondary endpoint was instrumental Activities of Daily Living (iADL). EXPEDITION3 did not demonstrate efficacy in the primary endpoint ADAS-cog14 but showed promising result in the secondary endpoint iADL. We use iADL for illustration.
Figure 3: Counterfactual uncertainty quantification can reduce the quantified uncertainty in efficacy estimation. The pie charts on the left show the proportion of total variance reduced by baselining while the pie charts on the right show the additional proportion reduced by counterfactual (instead of factual) calculations. The top two pie charts analyze data from the AbbVie study, while the bottom two pie charts analyze data from the EXPEDITION3 study. Overall, Counterfactual Uncertainty Quantification reduces 77.19% and 39.74% of the variances in the AbbVie study and the EXPEDITION3 study, respectively. The detailed calculation has been demonstrated in Appendix B.
5 A broader framework
A broader counterfactual estimation framework includes both Real Human Counterfactual Estimation (RHCE) and Digital Twin Counterfactual Estimation (DTCE). Our methodology is RHCE, baselining and self-controlling in the ETZ modeling principle with each human serving as their own control to reduce uncertainty quantification.
Another strategy for counterfactual estimation is to collect biomarker data on each patient, and train Digital Twin (DT) algorithms to predict each patient’s outcome from their biomarker values. In Digital Twin Counterfactual Estimation (DTCE), population $Rx:C$ efficacy is estimated by averaging across the patients the difference of their $Rx$ outcome and $C$ outcome.
Figure 4: Control-side and Equipoise counterfactual efficacy estimation.
Figure 5: Counterfactual estimation includes Real Human Counterfactual Estimation (RHCE) and Digital Twin Counterfactual Estimation (DTCE), either can be control-side or equipoise.
To explain potential biases that both RHCE and DTCE need to avoid, we separate Counterfactual estimation as either on the control-side or equipoise, as illustrated in Figure 4. Control-side counterfactual estimation, $CF_{control}$, takes only patients given $Rx$ and generates counterfactual $C$ outcomes. Equipoise counterfactual estimation, $CF_{equipoise}$, not only takes patients given $Rx$ and generates counterfactual $C$ outcomes but also generates counterfactual $Rx$ outcomes for patients given $C$. While $CF_{equipoise}$ has been standard practice, $CF_{control}$ is more recent thinking, in the DTCE setting.
In Section 5.2 that follows, we define traditional medicine versus modern targeted therapies, and explain how they have different response profiles. First, Figure 6 gives a visual preview that $CF_{control}$ may not even be applicable to traditional medicine, while $CF_{equipoise}$ however is unbiased for traditional medicine, but both $CF_{control}$ and $CF_{equipoise}$ can be biased for targeted therapies, as will be shown in Section 5.5.
5.1 Genesis of baseline covariate adjustment
The purpose of including baseline as a covariate in current MMRM analysis is to adjust for imbalance in the proportion of $Rx$ and $C$ treated patients in the data, to avoid bias. It is appropriate for traditional medicine with a baseline which is prognostic of outcome irrespective of the therapy received in the sense described on pages 19-22 of the FDA-NIH Biomarker Working Group (2016) BEST document.
If the data contains a substantial imbalance, then a real example in Hsu (1992) showed the marginal analysis implemented in the Means statement of SAS Proc GLM which does not adjust for the prognostic factor as a covariate can give very biased $Rx:C$ efficacy estimates.
Subsequently, starting with version 6.12 in 1996, SAS follows Chapter 7 of Hsu (1996) to provide unbiased multiple comparisons in the LSmeans statement, adjusting for covariate(s) presumed equally prognostic for $Rx$ and $C$ using Gauss-Markov least square means estimation to map the sample to a “balanced” population.
This practice has served the analysis of traditional medicine well for decades. However, targeted therapies are becoming quite common, and there is a possibility that for some of them baseline may be predictive of outcome in the sense described on pages 35-40 of the FDA/NIH (2016) BEST document, in which case one has to be more cautious.
5.2 From traditional medicine to targeted therapy
Traditional medicine
Traditional medicine is “one-drug-fits-all”, prescribing the same compound with no specific drug target to all patients for a particular indication. For example, label of olanzapine (Zyprexa®), which is indicated for patients with schizophrenia, states “The mechanism of action of olanzapine, as with other drugs having efficacy in schizophrenia, is unknown” (FDA (2009)).
Targeted therapies
Targeted therapies are personalized medicine, also called precision medicine, targeting specific disease pathways.
A monoclonal antibody, an amino acid chain (with compound suffix mab) that binds to a specific drug target is a targeted therapy. Immunotherapies for cancer such as nivolumab (Opdivo®) and pembrolizumab (Keytruda®) are examples of targeted therapies.
A tyrosine kinase inhibitor, a small molecule with compound suffix tinib, each intending to block a specific reaction within cells, is also a targeted therapy. Examples of tinibs are ibrutinib (for chronic lymphocytic leukemia CLL) and baricitinib (for rheumatoid arthritis RA).
Different patients receiving the same targeted therapy may derive differential benefits, because the extent to which the drug target presents itself in patients may differ by patient.
Figure 6: Conceptual plots of response profiles for both traditional medicine and targeted therapy. The left plot shows a traditional medicine has parallel profiles while the right plot shows a targeted therapy may well have non-parallel profiles. The Red dashed line in the top plot previews a situation where the apparent control-side counterfactual outcome profile is not parallel to the teal solid line, even for traditional medicine, in which case control-side counterfactual outcomes should not be produced, as will be explained in Section 5.4.
For both traditional medicine and targeted therapy, consider the simplest framework in which there are only two visits, baseline and milestone, and the true profiles are linear with random before treatment patient intercept $Z$ independent of random milestone measurement error. Referring to the patterns shown in Figure 6, we name the model used for traditional medicine as parallel traditional medicine model and name the model used for targeted therapy as non-parallel targeted therapy model.
A parallel traditional medicine model would be
$Y^{Rx[m]}_{r}$ $=$ $\prescript{{}_{Z}}{}{\theta}+\prescript{{}_{Z}}{}{\eta}^{Rx}+\prescript{{}_{Z}}{}{\beta}Z_{r}+\prescript{{}_{Z}}{}{\varepsilon}^{Rx[m]}_{r}$
$Y^{C[m]}_{s}$ $=$ $\prescript{{}_{Z}}{}{\theta}+\prescript{{}_{Z}}{}{\eta}^{C}+\prescript{{}_{Z}}{}{\beta}Z_{s}+\prescript{{}_{Z}}{}{\varepsilon}^{C[m]}_{s},$ (27)
where $\prescript{{}_{Z}}{}{\theta}$ is the expected outcome at baseline visit before either treatment is given, $\prescript{{}_{Z}}{}{\eta}^{Rx}$ is the fixed $Rx$ treatment effect, $\prescript{{}_{Z}}{}{\eta}^{C}$ is the fixed $C$ treatment effect, $\prescript{{}_{Z}}{}{\beta}$ is the same coefficient for $Z$ shared by treatment $Rx$ and $C$, $\prescript{{}_{Z}}{}{\varepsilon}^{Rx[m]}_{r}$ and $\prescript{{}_{Z}}{}{\varepsilon}^{C[m]}_{s}$ are the measurement error for $Rx$ and $C$ respectively, so patients sicker at the beginning remain relatively sicker at the end irrespective of $Rx$ or $C$.
For targeted therapy, a non-parallel targeted therapy model includes an additional interaction between $Z$ and treatment $Rx:C$
$Y^{Rx[m]}_{r}$ $=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{Rx}+\prescript{{}_{Z}}{}{\beta}^{Rx}Z_{r}+\prescript{{}_{Z}}{}{\epsilon}^{Rx[m]}_{r}$
$Y^{C[m]}_{s}$ $=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{C}+\prescript{{}_{Z}}{}{\beta}^{C}Z_{s}+\prescript{{}_{Z}}{}{\epsilon}^{C[m]}_{s},$ (28)
where $\prescript{{}_{Z}}{}{\alpha}$ is the expected outcome at baseline before either treatment is given, $\prescript{{}_{Z}}{}{\gamma}^{Rx}$ and $\prescript{{}_{Z}}{}{\gamma}^{C}$ are the fixed treatment effects for $Rx$ and $C$, $\prescript{{}_{Z}}{}{\beta}^{Rx}$ and $\prescript{{}_{Z}}{}{\beta}^{C}$ are different coefficients for $Z$ depending on whether the treatment is $Rx$ or $C$, $\prescript{{}_{Z}}{}{\epsilon}^{Rx[m]}_{r}$ and $\prescript{{}_{Z}}{}{\epsilon}^{C[m]}_{s}$ are the corresponding measurement errors, with patients sicker at the beginning benefiting more from targeted theory $Rx$ than from non-targeted therapy $C$ expressed as $\prescript{{}_{Z}}{}{\beta}^{Rx}$ being a steeper slope than $\prescript{{}_{Z}}{}{\beta}^{C}$.
Let $(B_{1},B_{2},\ldots,B_{p})$ be a vector of $p$ biomarkers and let $B=g(B_{1},B_{2},\ldots,B_{p})$ be a biomarker score calculated to be on the same scale as the response (the primary endpoint), with $B$ potentially predictive of outcome.
In practice, for patients $r$ and $s$, $B_{r}=Z_{r}+\prescript{{}_{B}}{}{E}^{[1]}_{r}$ and $B_{s}=Z_{s}+\prescript{{}_{B}}{}{E}^{[1]}_{s}$ with measurement errors are used instead of their (not observable) $Z$ values for prediction and ordinary regression is applied to the models (27) and (28). Burman et al. (2024) cites PROCOVA as an example of Digital Twin, deriving a prognostic score from historical patient-level baseline data and using it as a covariate in ANCOVA modeling to boost trial efficiency. Appendix D of Schuler et al. (2022) presents the predictors PROCOVA used to generate the prognostic score. Measurements of a number of the predictors, such as ADAS comprehension, CDR home and hobbies, and MMSE language, are subject to measurement error.
Then the parallel traditional medicine model becomes
$Y^{Rx[m]}_{r}$ $=$ $\prescript{{}_{B}}{}{\theta}+\prescript{{}_{B}}{}{\eta}^{Rx}+\prescript{{}_{B}}{}{\beta}B_{r}+\prescript{{}_{B}}{}{\varepsilon}^{Rx[m]}_{r}$
$Y^{C[m]}_{s}$ $=$ $\prescript{{}_{B}}{}{\theta}+\prescript{{}_{B}}{}{\eta}^{C}+\prescript{{}_{B}}{}{\beta}B_{s}+\prescript{{}_{B}}{}{\varepsilon}^{C[m]}_{s}$ (29)
while the non-parallel targeted therapy model becomes
$Y^{Rx[m]}_{r}$ $=$ $\prescript{{}_{B}}{}{\alpha}+\prescript{{}_{B}}{}{\gamma}^{Rx}+\prescript{{}_{B}}{}{\beta}^{Rx}B_{r}+\prescript{{}_{B}}{}{\epsilon}^{Rx[m]}_{r}$
$Y^{C[m]}_{s}$ $=$ $\prescript{{}_{B}}{}{\alpha}+\prescript{{}_{B}}{}{\gamma}^{C}+\prescript{{}_{B}}{}{\beta}^{C}B_{s}+\prescript{{}_{B}}{}{\epsilon}^{C[m]}_{s}$ (30)
as if $B_{r}$ and $B_{s}$ were measured without errors.
5.3 Noisy predictor causes attenuation
Regression modeling assumes predictors are measured without error, so what if some of $(B_{1},B_{2},\ldots,B_{p})$ have measurement error, and $B$ in turn has measurement error? For example, where as age and gender are presumably measured without error, baseline will have measurement error. Turns out if $B$ has measurement error, then estimated $Rx$ and $C$ effects will have so-called attenuation, and whether RHCE or DTCE is biased or unbiased for $Rx:C$ efficacy depends on how exactly that efficacy is defined, as will be shown in Section 5.4 and Section 5.5. We explain attenuation in the context that a biomarker score $B$ for each patient is used in RHCE, with $B$ being baseline.
Since the parallel traditional medicine model is a special case of non-parallel targeted therapy model in which $\prescript{{}_{Z}}{}{\beta}^{Rx}=\prescript{{}_{Z}}{}{\beta}^{C}$, we use non-parallel targeted therapy model to illustrate attenuation in general.
Model (28) which we take to be true have the predictor $Z$ measured without error, so
$\prescript{{}_{Z}}{}{\beta}^{Rx}=\frac{Cov(Y^{Rx[m]}_{r},Z_{r})}{Var({Z_{r}})}$ , $\prescript{{}_{Z}}{}{\beta}^{C}=\frac{Cov(Y^{C[m]}_{s},Z_{s})}{Var({Z_{s}})}.$
We show intuitively how attenuation occurs.
Substituting $Z_{r}=B_{r}-\prescript{{}_{B}}{}{E}^{[1]}_{r}$ and $Z_{s}=B_{s}-\prescript{{}_{B}}{}{E}^{[1]}_{s}$ into (28), we obtain
$Y^{Rx[m]}_{r}$ $=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{Rx}+\prescript{{}_{Z}}{}{\beta}^{Rx}(B_{r}-\prescript{{}_{B}}{}{E}^{[1]}_{r})+\prescript{{}_{Z}}{}{\epsilon}^{Rx[m]}_{r}$
$=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{Rx}+\prescript{{}_{Z}}{}{\beta}^{Rx}B_{r}+\prescript{*}{}{\epsilon}^{Rx[m]}_{r}$
$Y^{C[m]}_{s}$ $=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{C}+\prescript{{}_{Z}}{}{\beta}^{C}(B_{s}-\prescript{{}_{B}}{}{E}^{[1]}_{s})+\prescript{{}_{Z}}{}{\epsilon}^{C[m]}_{s}$
$=$ $\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{C}+\prescript{{}_{Z}}{}{\beta}^{C}B_{s}+\prescript{*}{}{\epsilon}^{C[m]}_{s}$
where error $\prescript{*}{}{\epsilon}^{Rx[m]}_{r}=\prescript{{}_{Z}}{}{\epsilon}^{Rx[m]}_{r}-\prescript{{}_{Z}}{}{\beta}^{Rx}(\prescript{{}_{B}}{}{E}^{[1]}_{r})$ and $\prescript{*}{}{\epsilon}^{C[m]}_{s}=\prescript{{}_{Z}}{}{\epsilon}^{C[m]}_{s}-\prescript{{}_{Z}}{}{\beta}^{C}(\prescript{{}_{B}}{}{E}^{[1]}_{s})$. The two errors $\prescript{*}{}{\epsilon}^{Rx}_{r}$ and $\prescript{*}{}{\epsilon}^{C}_{s}$ are correlated to the covariate $B_{r}$ and $B_{s}$ because they share the same component $\prescript{{}_{B}}{}{E}^{[1]}_{r}$ and $\prescript{{}_{B}}{}{E}^{[1]}_{s}$.
Take treatment $C$ as an example and similarly for $Rx$. This correlation between error and the covariate causes attenuation because with $Z_{s}$ and $\prescript{{}_{B}}{}{E}^{[1]}_{s}$ being independent,
$Cov(Y^{Rx[m]}_{s},B_{s})$ $=$ $Cov(Y^{Rx[m]}_{s},Z_{s}+\prescript{{}_{B}}{}{E}^{[1]}_{s})$
$=$ $Cov(Y^{Rx[m]}_{s},Z_{s})$
$Var(B_{s})$ $=$ $Var(Z_{s}+\prescript{{}_{B}}{}{E}_{s}^{[1]})$
$=$ $Var(Z_{s})+Var(\prescript{{}_{B}}{}{E}^{[1]}_{s})$
with an extraneous $Var(\prescript{{}_{B}}{}{E}^{[1]}_{s})$ for $Var(B_{s})$. Thus, comparing (30) with (28), we see $\prescript{{}_{B}}{}{\beta}^{C}$ is attenuated towards zero as emphasized in the measurement error literature (e.g., Carroll et al. (2006), p.43),
$\prescript{{}_{B}}{}{\beta}^{C}=\mathbbm{E}[\prescript{{}_{B}}{}{\hat{\beta}}^{C}]=\frac{Var(Z_{s})}{Var(Z_{s})+Var(\prescript{{}_{B}}{}{E}^{[1]}_{s})}\prescript{{}_{Z}}{}{\beta}^{C},$
similarly for $\prescript{{}_{B}}{}{\beta}^{Rx}$.
The attenuation can be quantified by the ETZ modeling principle in one special case where $B$ stands for baseline, the primary endpoint itself at first visit. In this case,
$B$ $=$ $Z+E^{[1]}$
$=$ $Z+E,~{}\mathrm{and}$
$\prescript{{}_{B}}{}{\beta}^{C}$ $=$ $\mathbbm{E}[\prescript{{}_{B}}{}{\hat{\beta}}^{C}]$ (31)
$=$ $\frac{Var(Z)}{Var(Z)+Var(E)}\prescript{{}_{Z}}{}{\beta}^{C},$
noting the extra $Var(E)$.
We show below the presence of measurement error $\prescript{{}_{B}}{}{E}$ or $E$ in using baseline or any other biomarker $B$ score as a surrogate for the true patient effect $Z$ causes attenuation in estimating the regression coefficient of $B$, which in turn can cause bias in estimating $Rx:C$ efficacy. While future ETZ research might mitigate this bias, for now we show that estimated counterfactual efficacy is unbiased provided that efficacy is defined as “on average”.
5.4 Possible bias caused by attenuation
Bias would occur if one predicts efficacy in subgroups of patients (e.g., Alzheimer’s disease patients with mild or severe symptoms at baseline) when RHCE is modeled (without stratification) for the entire population or DTCE is trained on the entire population.
These biases can be visualized. We use the outcome instrumental Abilities of Daily Living (iADL) for Alzheimer’s disease to illustrate. Abilities to manage finances and to prepare food are some of the items measured by iADL.
We simulate potential $Rx$ and $C$ outcomes at milestone visit for 500 patients in a counterfactual setting from the ETZ modeling principle with first visit, baseline (before randomization) measurements
$Y^{[1]}_{i}$ $=$ $Y^{[1]}_{i}(Rx)$
$=$ $Y^{[1]}_{i}(C)$
$=$ $Z_{i}+E^{[1]}_{i},\ i=1,\dots,500$
and milestone visit measurements
$Y^{[m]}_{i}(Rx)$ $=$ $Z_{i}+Traj^{Rx[m]}_{i}+E^{Rx[m]}_{i},$ (32)
$Y^{[m]}_{i}(C)$ $=$ $Z_{i}+Traj^{C[m]}_{i}+E^{C[m]}_{i},$
$i=1,\dots,500.$
In ETZ modeling, the three components $Z$, $Traj$, and $E$ follow different Normal, independent, distributions. $Z_{i}$ is generated from a Normal distribution with mean $\alpha$ and variance $Var(Z)$. $Traj^{Rx[m]}_{i}$ and $Traj^{C[m]}_{i}$ are generated from Normal distributions with means $\mu^{Rx}$ and $\mu^{C}$ and common variance $Var(Traj)$, independently. $E^{[1]}_{i}$, $E^{Rx[m]}_{i}$ and $E^{C[m]}_{i}$ are generated from Normal distributions with mean 0 and variance $Var(E)$, independently. In the counterfactual setting, outcomes $Rx$ and $C$ for patient $i$ share the same $Z_{i}$.
Summary statistics from the EXPEDITION3 study and ETZ modeling principle in Table 4 are used for data generation.
Variance SD
Visit 1 (baseline, $Y^{[1]}$) 64.58 8.04
Visit 7 (milestone visit, $Y^{[m]}$) 135.39 11.64
Change-from-baseline ($Y^{[change]}$) 92.37 9.61
Intercept ($Z$) 53.80 7.34
Trajectory ($Traj$) 70.81 8.42
Measurement Error ($E$) 10.78 3.28
Table 4: Variance and standard deviation (SD) of iADL at Visits 1 and $m$, and change-from-baseline of the EXPEDITION3 study are as reported in Honig et al. (2018). Vriances of the Intercept, Trajectory and Measurement Error are computed by the ETZ transformation.
Figure 7: This plot shows attenuation when $B$ is a surrogate of the true $Z$ measured with error. The slope regressing $Y$ on $B$ tends to be closer to zero, compared to the true slope regressing $Y$ on $Z$. Black arrows in the plot show attenuation bias in estimated $C$ effect, due to measurement error in biomarker score $B$ as a surrogate of the true $Z$. Treatments are indicated by color while shape differentiates $B$ from $Z$. Red represents control $C$ while teal represents treatment $Rx$. Red round points are $C$ versus $B$ measurements, with the red dashed line being the fitted $C$ versus $B$ regression line. red dashed line are $C$ versus $Z$ measurements, with the red solid line being the fitted $C$ versus $Z$ regression line. Teal triangle points are $Rx$ versus $Z$ measurements, with the teal solid line being the fitted $Rx$ versus $Z$ regression line.
Figure 8: This plot shows how bias in $Rx$ versus $C$ efficacy estimation depends on the target value of outcome $C$. The orange points represent outcome $Rx$ (vertical axis) as predicted by true $Z$ plotted against outcome $C$ (horizontal axis) as predicted by true $Z$, with the orange line being the prediction line. The blue points represent outcome $Rx$ (vertical axis) as predicted by true $Z$ plotted against outcome $C$ (horizontal axis) as predicted by $B(=Z+E)$, with the blue line being their fitted line. $Rx$ versus $C$ efficacy, the average difference of response given $Rx$ and response given $C$, is biased upward when target $C$ value is below average (with the blue points trending above the orange points), downward when target $C$ value is above average (with blue points trending below orange points).
Using baseline or, more generally, a biomarker score $B$ to predict the outcome is tempting because $Z$ is not observable. However, as Figures 7 and 8 show, one should be aware that predicting the expected outcome at a $B$ value off-center may incur bias if $B$ has measurement error.
5.5 Estimated average $Rx{:}C$ efficacy is unbiased
Population effect of treatment $Rx$ or $C$ is defined as response $Y$ averaged over infinitely many patients, each with their value $Z$. In the population space, what is of interest is the value $\mathbbm{E}[Y]$ at $\mathbbm{E}[Z]$.
One intuitive way of seeing why the estimated average effect is unbiased is to imagine regressing $Y$ on $Z$ in (28) versus regressing $Y$ on $B$ in (30), with a rather large number of patients. Attenuation, which occurs when $B$ is a surrogate of $Z$ but measured with error, means the slope regressing $Y$ on $B$ in (30) tends to be closer to zero, compared to the true slope regressing $Y$ on $Z$ in (28).
A familiar experience in simple linear regression is the fitted regression line goes through the central point $(\bar{X},\bar{Y})$ of average predictor and response values. Thus, imagining the clinical study being repeated, one can visualize the pairs of regression lines from fitting (28) and (30) pivoting/rotating around the same central point, as depicted in Figure 7. Since $\mathbbm{E}[B]=\mathbbm{E}[Z]$, even though the two regression lines have different slopes, their values at $\mathbbm{E}[Z]$ (average value of $Z$) are the same.
To explain unbiasedness mathematically, we define efficacy more generally as measured at $\mathbbm{E}[Z]\pm c$ where $c$ is any non-negative constant. (Population efficacy as described earlier would be at $\mathbbm{E}[Z]$ with $c=0$.) We prove that:
  • For traditional medicine, estimated equipoise $Rx:C$ efficacy defined at $\mathbbm{E}[Z]\pm c$ is unbiased for all $c\geq 0$.
  • For targeted therapy, control-side and equipoise $Rx:C$ efficacy estimates are only unbiased when $c=0$, but can be biased when efficacy $\mathbbm{E}[Z]\pm c$ is defined off-centered with $c>0$.
Traditional medicine has parallel profiles. First, we caution control-side $Rx:C$ efficacy estimation is biased and should not be done, for the following reason. For traditional medicine, treatment $Rx$ and treatment $C$ share the same coefficient for $baseline$ in both model (27) and (29). For control-side estimation, counterfactual $C$ outcomes are predicted through model (29) with attenuated baseline coefficient $\prescript{{}_{B}}{}{\beta}=\frac{Var(Z)}{Var(Z)+Var(\prescript{{}_{B}}{}{E})}\prescript{{}_{Z}}{}{\beta}$, while factual $Rx$ outcomes follow the true model (27) with coefficient of baseline, $\prescript{{}_{Z}}{}{\beta}$. Therefore, as shown in Figure 6, control-side estimation makes the apparent profiles not parallel for traditional medicine and should not be done.
Now we prove equipoise $Rx:C$ efficacy estimation is unbiased for traditional medicine mathematically, no matter what value $c\geq 0$ efficacy $\mathbbm{E}[Z]\pm c$ is defined at. For equipoise counterfactual estimation, efficacy is measured within each of $Rx$ and $C$ first and then combined. Efficacy is measured as the difference between a factual measured outcome and its corresponding counterfactual outcome within each treatment arm which, assuming the numbers of patients in $Rx$ and $C$ are equal, is
$\frac{1}{2}[(\prescript{{}_{Z}}{}{\theta}+\prescript{{}_{Z}}{}{\eta}^{Rx}+\prescript{{}_{Z}}{}{\beta}(\mathbbm{E}[Z]\pm c))$
$-(\prescript{{}_{B}}{}{\theta}+\prescript{{}_{B}}{}{\eta}^{C}+\prescript{{}_{B}}{}{\beta}(\mathbbm{E}[Z]\pm c))]$
$+\frac{1}{2}[(\prescript{{}_{B}}{}{\theta}+\prescript{{}_{B}}{}{\eta}^{Rx}+\prescript{{}_{B}}{}{\beta}(\mathbbm{E}[Z]\pm c))$
$-(\prescript{{}_{Z}}{}{\theta}+\prescript{{}_{Z}}{}{\eta}^{C}+\prescript{{}_{Z}}{}{\beta}(\mathbbm{E}[Z]\pm c))]$
$=\mathbbm{E}[Y^{Rx[m]}_{r}]-\mathbbm{E}[Y^{C[m]}_{s}].$
Hence, equipoise counterfactual estimation is unbiased for traditional medicine.
We discuss both control-side and equipoise counterfactual estimations for targeted therapy. For control-side counterfactual estimation of targeted therapy, efficacy is measured as the difference between the factual outcome of a patient receiving $Rx$ and the corresponding counterfactual $C$ outcome:
$(\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{Rx}+\prescript{{}_{Z}}{}{\beta}^{Rx}(\mathbbm{E}[Z]\pm c))-$
$(\prescript{{}_{B}}{}{\alpha}+\prescript{{}_{B}}{}{\gamma}^{C}+\prescript{{}_{B}}{}{\beta}^{C}(\mathbbm{E}[Z]\pm c))$
$=$ $(\mathbbm{E}[Y^{Rx[m]}_{r}]-\mathbbm{E}[Y^{C[m]}_{s}])\pm$
$c(\prescript{{}_{Z}}{}{\beta}^{Rx}-\frac{Var(Z)}{Var(Z)+Var(\prescript{{}_{B}}{}{E})}\prescript{{}_{Z}}{}{\beta}^{C})$
so there is bias if $c>0$.
Equipoise counterfactual estimation for efficacy in targeted therapy can be expressed as:
$\frac{1}{2}[(\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{Rx}+\prescript{{}_{Z}}{}{\beta}^{Rx}(\mathbbm{E}[Z]\pm c))$
$-(\prescript{{}_{B}}{}{\alpha}+\prescript{{}_{B}}{}{\gamma}^{C}+\prescript{{}_{B}}{}{\beta}^{C}(\mathbbm{E}[Z]\pm c))]$
$+\frac{1}{2}[(\prescript{{}_{B}}{}{\alpha}+\prescript{{}_{B}}{}{\gamma}^{Rx}+\prescript{{}_{B}}{}{\beta}^{Rx}(\mathbbm{E}[Z]\pm c))$
$-(\prescript{{}_{Z}}{}{\alpha}+\prescript{{}_{Z}}{}{\gamma}^{C}+\prescript{{}_{Z}}{}{\beta}^{C}(\mathbbm{E}[Z]\pm c))]$
$=(\mathbbm{E}[Y^{Rx[m]}_{r}]-\mathbbm{E}[Y^{C[m]}_{s}])$
±$\pm\frac{c}{2}(\prescript{{}_{Z}}{}{\beta}^{Rx}-\frac{Var(Z)}{Var(Z)+Var(\prescript{{}_{B}}{}{E})}\prescript{{}_{Z}}{}{\beta}^{C})$
±$\pm\frac{c}{2}(\frac{Var(Z)}{Var(Z)+Var(\prescript{{}_{B}}{}{E})}\prescript{{}_{Z}}{}{\beta}^{Rx}-\prescript{{}_{Z}}{}{\beta}^{C})$
so there is bias if $c>0$. That would happen if the target patient population in the actual RCT does not align with the patient population in the learning study. Table 5 advises extraordinary caution in this situation using two examples.
MMSE scoring
Study Mild/Early Moderate
EXPEDITION1 $\checkmark$ $\checkmark$
Start of EXPEDITION2 $\checkmark$ $\checkmark$
End of EXPEDITION2 $\checkmark$
EXPEDITION3 $\checkmark$
Training ProCovA $\checkmark$
Applying ProCovA $\checkmark$ $\checkmark$
Table 5: From Section 5.4, and experience from EXPEDITION1/2/3 (which did not meet their primary objectives), extraordinary caution is advisable in predicting potential outcome when the predictor contains measurement error and the learning study patient population differs from the application study patient population. EXPEDITIONS 1, 2, and 3 staged Alzheimer’s disease patients by the Mini–Mental State Examination (MMSE) score, with an MMSE score of 20 to 26 classified as mild, while a score of 16 to 19 is moderate. From Honig et al. (2018), MMSE has an estimated $Var(E)$ of 2.06 which is not ignorable since the estimated MMSE difference between solanezumab and the placebo is 0.49. EXPEDITION 2 started by targeting mild and moderate patients, but changed to targeting mild patients after learning from EXPEDITION1’s result. EXPEDITION3 targeted only mild patients after learning from data on mild patients pooled from EXPEDITION1 and EXPEDITION2. From Schuler et al. (2022), it appears ProCovA was applied to mild and moderate patients with MMSE between 14 and 26 in the DHA supplement study, while the algorithm appears to have been trained on early stage patients.
6 Guidelines for applying the ETZ modeling principle
ETZ is a modeling principle that enables Counterfactual Uncertainty Quantification (CUQ) for randomized studies with Before-and-After Repeated Measures (BAtRM). Scientists faced with highly variable outcome measure (i.e., large measurement error) situations can benefit from this principle, provided that they can extract the three inputs needed by ETZ, working out details for proper CUQ in their settings themselves.
For example, highly variable responses to therapies for amyotrophic lateral sclerosis (ALS, affecting $\approx$ 31,000 patients in the U.S.), measured on the ALS Functional Rating Scale–Revised (ALSFRS-R) scale (FDA (2021)) can be modeled with a random coefficient BAtRM model (e.g., using the RANDOM statement in Proc Mixed of SAS), providing inputs needed by ETZ. As another example, RCTs in immunology such as for systemic lupus erythematosus (SLE, affecting $\approx$ 200,000 patients in the U.S.), often are BAtRM studies with binary endpoints. For binary and time-to-event outcomes, efficacy is often measured as a ratio. The CUQ principle for reducing uncertainty quantification can be applied to ratio efficacy, provided one is aware of the following:
  1. 1.
    Mixing not-logic-respecting efficacy measures Odds Ratio (OR) and Hazard Ratio (HR) in subgroups attenuates them, even without measurement error in (biomarker) predictors, see Liu et al. (2022);
  2. 2.
    Computer packages currently give misleading stratified analysis results for all ratio efficacy (logic-respecting or not), because the prognostic effect is ignored, see Liu et al. (2023).
However, following the Subgroup Mixable Estimation (SME) principle, correct causal inference for binary and time-to-event outcomes can be obtained for logic-respecting rato efficacy such as Relative Response (RR) and Ratio-of-Time, see Ding et al. (2016), Lin et al. (2019), the discussion paper Liu et al. (2022), and its rejoinder Liu et al. (2022).
7 Key messages
ETZ is an uncertainty quantification (UQ) principle for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures (BAtRM). Conducting RCT with BAtRM is standard practice in therapeutic areas affecting more than half a billion patients worldwide. By modeling individual patient outcomes in BAtRM studies, ETZ enables counterfactual uncertainty quantification (CUQ). CUQ is typically smaller than current factual uncertainty quantification. Statistical or digital counterfactual point estimation involving a biomarker with measurement error is unbiased for efficacy in the entire population but biased in subgroups.
8 Acknowledgments
The ETZ modeling thinking started with Haiyan Xu saying “Dr. Hsu, you only need three numbers!”
We thank the AE for on-point comments, particularly for requesting us to comment on ProCovA.
9 Conflicts of interest
X.W., Y.H., S-Y.T. declare no competing interests.
Y. L. is an employee and shareholder of Eli Lilly.
J.C.H. is a professor emeritus of the Ohio State University and a consultant to Eli Lilly.
Appendix A Glossary of terminologies
BAtRM: Before-and-After treatment Repeated Measures. Effect of a treatment: Expected outcome given treatment $Rx$ or $C$. Efficacy: Expected differential outcome between treatments $Rx$ and $C$. Attenuation: Directional bias (within each treatment arm) in estimating slope of the treatment effect when predictor contains measurement error. Factual study: A study in which each patient is given either $Rx$ or $C$. Counterfactual study: A study in which each patient is given both $Rx$ and $C$. Counterfactual efficacy: Expected differential between $Rx$ and $C$ outcomes of patients in a counterfactual study. Counterfactual estimation (CE): Unbiased point estimation of counterfactual efficacy from a factual study. Counterfactual uncertainty quantification (CUQ): Quantification of uncertainty (such as variance) of a counterfactual point estimate calculated as if it were from a counterfactual study. Targeted therapies: Therapies such as tyrosine-kinase inhibitors and monoclonal antibodies that target specific pathways, constituting a large portion of “personalized” or “precision” medicine. Real Human Counterfactual Estimation (RHCE): Counterfactual point estimation (CE) with Counterfactual Uncertainty Quantification (CUQ) using each real human patient as self-control. Digital Twins Counterfactual Estimation (DTCE): Counterfactual point estimation (CE) using biomarkers to predict for each $Rx$ patient its digital $C$ twin’s outcome and vice versa, then averaging the outcome differentials.
Appendix B Details of CUQ variance reduction calculations
Detailed calculations of the two-step CUQ process that led to the pie chart for EXPEDITION3 in Figure 3 are given here. Calculations for the AbbVie data are similar.
Table 4 presents summary statistics for EXPEDITION3 along with their ETZ variance components. Detailed calculation showing baselining reduces variance in milestone visit by 43.02 which is 31.77% of total variance is as follows.
$Var(Y^{[m]})$ $=$ $Var(Z)+Var(Trt)+Var(E)$
$=$ 135.39135.39 135.39135.39
$Var(Y^{[change]})$ $=$ $Var(Trt)+2\times Var(E)$
$=$ 92.3792.37 92.3792.37
$Var(Y^{[m]})-Var(Y^{[change]})$ $=$ $Var(Z)-Var(E)$
$=$ $53.802-10.778$
$=$ 43.0243.02 43.0243.02
$\frac{Var(Y^{[m]})-Var(Y^{[change]})}{Var(Y^{[m]})}$ $=$ $31.77\%.$
Detailed calculation showing the counterfactual aspect of CUQ further reduces variance by 21.56 beyond baselining by 11.67% is as follows.
$Var(Y^{f})$ $=$ $2\times Var(Trt)+4\times Var(E)$
$=$ $2\times 70.809+4\times 10.778$
$=$ 184.73184.73 184.73184.73
$Var(Y^{cf})$ $=$ $2\times Var(Trt)+2\times Var(E)$
$=$ $2\times 70.809+2\times 10.778$
$=$ 163.174163.174 163.174163.174
$Var(Y^{f})-Var(Y^{cf})$ $=$ $2\times Var(E)$
$=$ $184.73-163.174$
$=$ 21.55621.556 21.55621.556
$\frac{Var(Y^{f})-Var(Y^{cf})}{Var(Y^{f})}$ $=$ $11.67\%.$
Hence, CUQ reduces
$31.77\%+11.67\%\times(1-31.77\%)=39.73\%$ (34)
of the total variance in EXPEDITION3.